When pO2 = 4*p50 of myoglobin, the fractional saturation (Y O2) is about?
Answer #1

Formula: YO2 = pO2/KD + pO2

Here, KD = p50

According to the given data, p50 = pO2/4

Therefore, YO2 = pO2/(pO2/4 + pO2)

i.e. YO2 = 0.8